# FAST functional connectivity implicates P300 connectivity in working memory deficits in Alzheimer’s disease

### v1

@@ II-C Network analysis of FAST connectivity
The clustering coefficient quantifies the number of connected triangles in a network and, thus, the tendency of nodes to cluster together. This weighted version multiplies the triangle weights together, with larger values where all triangle weights are large. The average value for each temporal window emphasises the strongly clustered components in the signal. The computation is fairly straightforward with the sum of the main diagonal of the cube of the tensor divided by the number of nodes (EEG electrodes). We can make two interesting observations here, taking into account the Law of Large numbers stating that as the number of independent samples increases the empirical mean converges to the true mean we can conclude that increasing the number of electrodes and thus nodes (and thus independent samples) will give us a more stable and reliable estimate of the mean for a given time-step allowing for a greater temporal resolution, this has been implicated in previous studies, where there is strong evidence showing that as a result of reducing electrode density networks tended to get skewed [16], this effect was most prominent below 64 electrodes. Also, the Central Limit Theorem tells us that the estimate will converge to a normal distribution as the number of independent samples increases, suggesting that increasing the number of electrodes would allow the network metric estimates to follow a more Gaussian distribution, providing better estimates of the mean to allow for more robust statistical testing.
II-D Simulations
@@
We utilized open-source MATLAB functions provided by Yeung et al.[17, 18] to generate the simulated EEG data. The simulated data consists of two key components: noise and signal. The noise component is generated to mimic the power spectrum of a typical human EEG recording. The signal component is parameterized to describe the position of the centre of the peak, its frequency, and its amplitude. These parameters enable us to create sample ERPs, which serve as the basis for testing the effectiveness of our method. The EEG simulation functions provide a setup with 31 electrodes, each sampled at a frequency of 200Hz, with an epoch duration of 0.8 seconds. To generate independent samples, we averaged the random signals over varying numbers of trials, resulting in single 31x200-dimensional samples that closely resemble real EEG data.
@@
For our experimental setup, we aimed to replicate conditions akin to typical comparisons between participants in time-locked Visual Short-Term Memory (VSTM) tasks. We created 20 independent samples consisting solely of EEG noise, aligning with the EEG power spectrum of a typical human. In parallel, we generated 20 independent samples with specific ERPs, including the N100 and P300 components. The N100 component was parameterized with an amplitude of -5 (typically negative) and a frequency of 5 Hz, with a centre position at 25 frames (around 100ms) considering the 0.8-second epoch. The P300 component was parameterized with an amplitude of 5 (positive and larger than N100) and a frequency of 5 Hz, with a centre position at 75 frames (around 300ms) within the 0.8-second epoch. To introduce realistic variability, the functions incorporate temporal jitter at the onset of the ERPs, mirroring the kind of activity observed in actual EEG ERP data. We then added random Gaussian white noise to the samples containing the simulated ERPs to test the robustness of our FAST Filtering method in the presence of variable levels of external noise. This approach allows us to rigorously test the performance of our method in distinguishing between the presence and absence of these specific ERP components in simulated EEG signals.
@@
II-E Visual short-term memory data
We use the data from Pietto et al. [9] with 10 patients having Alzheimer’s due to a familial gene, which is a younger age group with Mild Cognitive Impairment (MCI) compared to 10 age and education-matched Control patients. This is a 64-electrode setup with a sampling frequency of 500Hz over a one-second epoch of binding or shape tasks. We have a second data set of Sporadic AD patients with 13 Patients with MCI due to AD due to no specific genetic cause and 17 age and education-matched controls. This is a 128-electrode setup with a sampling frequency of 255Hz with the same VSTM tasks over a one-second epoch. For full details of these data sets the reader is referred to [9]. Both datasets consist of MCI patients at risk of AD and Controls performing a visual short-term memory task that assesses, via two conditions, memory for shapes and for shape-colour binding.
@@ II-E Visual short-term memory data
Signal pre-processing was performed to get signals band-passed into Delta (0.01-4Hz), Theta (4-8Hz), Alpha (8-12Hz), Beta (12-30Hz) and Gamma ($>$30Hz) frequencies.
@@
The network analysis was done using FAST Connectivity. We first computed the long-term connectivity over the whole epoch for each patient and control in each task using Pearson’s correlation coefficient. We then applied Definition 1 to create our FAST filter of overall general VSTM task activity. A separate filter was created for the familial AD dataset and for the Sporadic AD dataset due to them being different experiments with different numbers and positioning of electrodes and different sampling frequencies. We then split the FAST Connectivity tensors into 10 0.1 second, non-overlapping temporal windows by averaging over smaller time windows to give us 10 matrices of FAST connectivity for each participant with a high temporal resolution.
II-F Statistical Methods
@@
We designate as a result of clinical interest as a given temporal window where there is a significant difference between patients and controls in the binding task and no significant difference in the shape task as this points to a specific binding deficit in AD. Non-parametric Wilcoxon rank-sum test are performed to assess for statistical significance between patients and controls. These are computed at each 0.1s temporal window between the vectors of mean network metrics for patients and controls at each temporal window. This is repeated for the mean edge weights and the mean weighted clustering coefficient values.The direction and size of the differences are calculated using Cohen’s d effect size. In our experiments, a negative value would indicate a greater connectivity in the AD patients for the given VSTM task.
@@
In order to account for the multiple comparisons we applied the Benjamini-Hochberg False Discovery Rate (FDR) correction to account for multiple temporal significance testing. This was done at the 10 and 5 percent level. While 5 percent is often held as the strict standard, the 10 percent level allows us to look for sensitivity pointing towards reproducibility across datasets– i.e. where one dataset passed FDR at a given timepoint at 5 percent and the other at 10 percent.
III Results
III-A FAST Connectivity Simulation
@@
While there is a theoretical rationale to support our novel methodology, it is not a given that this follows through meaningfully in practice. Therefore, a simulation in a general case scenario is implemented to show the effectiveness of our method at picking up relevant Event-Related Potentials. This is tested across varying noise levels and the number of task trials. We performed a three-way comparison between the unfiltered node-function, individual GVD filters and our FAST filter.
The ERPs we chose to replicate are the N100 and P300. The N100 is characterized as a negative deflection in the EEG signal that occurs around 100 milliseconds after cognitive stimuli onset. The N100 typically peaks at around 100 milliseconds after onset and is usually attributed to early sensory processing in response to various stimuli. The N100 has been previously implicated in various neurological disorders such as Schizophrenia, Attention Deficit Hyperactivity Disorder (ADHD) and even Alzheimer’s Disease. The P300 is characterized by a positive deflection in the EEG signal and usually occurs around 300 milliseconds post the presentation of stimuli. The P300 can be influenced by the given task the participant is involved in and is associated with the evaluation of the relevance of stimuli. The P300 has been heavily researched in AD [19, 20] and is associated with decision-making and working memory. We simulate these characteristics using the EEG model of Yeung et al.[17, 18].
Our aim is to evaluate the effectiveness of our methodology in detecting these simulated ERPs compared to a ‘control’ group where the ERPs are not present. We first use a Matrix of 1’s, the same dimensions as our graph-variate signal, with 0’s on the main diagonal to represent the unfiltered node function. We used the squared difference as our node function. Table 1 shows the log of the Wilcoxon rank-sum test $p$-values between the two groups of simulated participants at the time steps at 1000 trials with a very high level of external noise. Benjamini-Hochberg FDR correction is applied to account for multiple comparisons.
TABLE I: P-values over High External Noise Levels and 1000 Trials
Network Metric FAST Filtered Unfiltered
Clustering Coefficient $2.96\times 10^{-6}$ 0.6344
Edge Weight $1.699\times 10^{-7}$ 1
The goal of filtration with long-term stable connectivity in Graph Variate Signal Analysis is to improve the ability to detect important changes in activity in the presence of large amounts of noise. In order to show the benefit of the FAST filter compared to the unfiltered node function we added large amounts of white noise to the participants with the ERPs, essentially ‘covering up’ the ERPs to reach the point at which the unweighted node function can no longer efficiently pick out the simulated P300 and N100 ERPs but the FAST Filter can. Random Gaussian white noise is added randomly to each of the electrodes equally in the simulated setup (Note, we are adding random white noise which is distinct from the signal generated by the MATLAB functions that resemble the power spectrum of a typical human EEG recording). Figure 1 shows the $p$-values detected by the unfiltered modulus of the squared difference at the pre-determined P300 timestep. It is clear that it is unable to pick up any significant changes at this time point due to the external noise. The FAST filter however, picks up the significant differences at the time step with both the mean edge weights and clustering coefficient.
It is now evident that the unfiltered node function is failing to pick up the ERPs above a certain noise level limit. When we compared using individual filters for each participant, we found ‘significant differences’; however, on further inspection, Figure 2 shows that the external noise results in the filters failing and every time-step is ’significantly different’, whereas the global filter maintains the non-different and different time steps between both signal groups. Essentially the underlying stable support matrix is influencing the dynamic connectivity too much and thus detecting ’significant differences’ in temporal windows when there are none simply due to differences in the filters. The unbiased nature of the global filter in FAST connectivity helps overcome this issue.
Figure 1: Individual filters (Top) vs FAST Filters (Bottom), Blue for Clustering Coefficient,Green for Edge Weights
We found that, with the FAST filter, while sometimes a network metric might not pick out a significant result, in our experiments it never picked out spurious results as significant. I.e the false positive rate also appears to be very low. We have shown that the FAST filter outperforms the individual filters and the unweighted node function in a general simulated experimental setup similar to time-locked VSTM task comparisons.
@@
We decided to test this more rigorously by varying levels of added Gaussian noise at different numbers of trials for each simulated participant EEG. The figure below shows the results of this as a heat-map matrix. We extracted the FDR corrected $p$-values at the P300 ERP time steps of interest (pre-determined to exist at these times-steps) thus we could compare the ability of the unfiltered and FAST filtered methods to pick out significant differences at these points. Lighter colours indicate more significant $p$-values while black represents a non-significant result. We tried trials ranging from 50-300 corresponding to typical real-life experiments where EEG signals are recorded.
@@
Figure 2: The $p$-values at a pre-determined time-step of the simulated P300 at increasing levels of trial size and external Gaussian noise. FAST Filtered (Bottom) vs Unfiltered (Top).
We can see the FAST filter’s robustness to increasing levels of noise with the edge weights picking up significant results at the correct time steps in almost all cases in the 100-300 trial range. At 50 trials, it still performs relatively well but there is a decreasing performance as more noise is added. The trend shows that as the number of trials increases the detection ability improves while increasing noise decreases this detection ability. The mean edge weights in the unfiltered case fail to detect these simulated ERPs in almost all levels of noise and trial sizes. Overall, the simulations are promising showing the clear benefit of the FAST filter compared to the unfiltered case and its evident robustness to noise.
@@
Another important consideration in analyzing Dynamic functional connectivity in small temporal windows is the temporal resolution we can achieve while still maintaining robustness to noise. Previous methods depended heavily on the length of the sliding window in finding a trade-off between temporal resolution and robustness to noise. We decided to test the window length dependency of FAST connectivity by setting the number of windows equal to the sampling rate, i.e. maximum temporal resolution. Figure 3 shows the results of this at varying noise levels and trial sizes.
Figure 3: Performance at maximum temporal resolution of FAST Filtered simulated EEG functional connectivity
@@
The findings exhibit promise, revealing that despite a slight reduction in performance when utilizing maximum temporal resolution with 200 windows in contrast to 10 temporal windows, the decline is minimal and potentially insignificant, particularly with an optimal number of trials. This observation underscores the robustness of FAST connectivity in variance to window length variations, highlighting its comparative advantage over existing methodologies in capturing dynamic functional connectivity changes. Notably, at 200 trials, FAST connectivity demonstrates consistent detection across all noise levels. We note that at 50 trials, FAST connectivity can detect meaningful changes in functional connectivity; working memory tasks often require participants to engage in sustained cognitive effort, leading to potential cognitive fatigue, especially in prolonged experimental sessions. FAST connectivity shows potential to address this challenge by exploring the feasibility of achieving accurate ERP detection with a lower number of trials. Additionally, we should take into account economic considerations prevalent in low-income countries, where optimizing experimental protocols can significantly reduce costs associated with data acquisition and analysis.
@@
III-B Application to visual short-term memory binding in Alzheimer’s disease

### v2

@@ 2.3 Network analysis of FAST connectivity
[1]
We limited our choice to these two network metrics as we did not perform any binarization on the connectivity profiles in order to maximize the information we can gain. Thus, we are just analyzing changing edge weights of the completely connected graph, as a consequence of this the weighted clustering coefficient and the edge weights represented the general topological distribution of the connectivity profile effectively and provided sufficient results.
Furthermore, at this initial stage of our methodological development, our goal is to identify significant changes in global functional connectivity. To achieve this, we are not conducting a node-by-node analysis. This approach is crucial for small sample-sized datasets, as nodal-level testing greatly reduces statistical power due to multiple comparison corrections. Instead, we are adopting a hypothesis-free approach regarding specific brain regions of interest.
The mean of the edge weights in this scenario essentially captures the Dirichlet energy of the entire instantaneous filtered connectivity profile thus provides a reliable measure of instantaneous connectivity as indicated in previous studies[2].
The clustering coefficient, on the other hand, quantifies the number of connected triangles in a network and, thus, the tendency of nodes to cluster together. This weighted version multiplies the triangle weights together, with larger values where all triangle weights are large. The average value for each temporal window emphasises the strongly clustered components in the signal. The computation is fairly straightforward with the sum of the main diagonal of the cube of the tensor divided by the number of nodes (EEG electrodes).
@@
We can make two interesting observations here, taking into account the Law of Large numbers stating that as the number of independent samples increases the empirical mean converges to the true mean we can conclude that increasing the number of electrodes and thus nodes (and thus independent samples) will give us a more stable and reliable estimate of the mean for a given time-step allowing for a greater temporal resolution, this has been implicated in previous studies, where there is strong evidence showing that as a result of reducing electrode density networks tended to get skewed [29], this effect was most prominent below 64 electrodes. Also, the Central Limit Theorem tells us that the estimate will converge to a normal distribution as the number of independent samples increases, suggesting that increasing the number of electrodes would allow the network metric estimates to follow a more Gaussian distribution, allowing us to exploit the various statistical approaches that assume Gaussianity.
2.4 Wavelet Power Spectra Analysis
The Wavelet Transform [36] is a powerful tool for analyzing temporally varying signal data. It allows for the decomposition of a signal into components localized in both time and frequency domains. One of the commonly used wavelets for continuous wavelet transform (CWT) is the Morlet wavelet, which is particularly useful for detecting oscillatory patterns in the signal.
The continuous wavelet transform of a signal $x(t)$ using a mother wavelet $\psi(t)$ is defined as:
[…]
where $a$ is the scale parameter, which controls the dilation of the wavelet, $b$ is the translation parameter, which controls the translation of the wavelet, and $\psi^{*}(t)$ is the complex conjugate of the mother wavelet $\psi(t)$.
The Morlet wavelet is defined as:
$\psi(t)=\pi^{-\frac{1}{4}}e^{j\omega_{0}t}e^{-\frac{t^{2}}{2}}$ (14)
where $\omega_{0}$ is the central frequency of the wavelet.
The wavelet transform can be viewed as a convolution of the signal $x(t)$ with a set of wavelet functions. Each wavelet function is a scaled and shifted version of the mother wavelet $\psi(t)$. The wavelet coefficients $W_{x}(a,b)$ represent the correlation between the signal and the wavelet at different scales $a$ and positions $b$.
To analyze the EEG signals, we split the data into ten disjoint temporal windows. For each window, we computed the power spectrum for each channel using the wavelet coefficients. This allowed us to analyze the signal in the frequency-time domain. This process helps in understanding how the power of different frequency components of the signal varies over time. In this case we used the Morlet wavelet as our mother wavelet of choice.
2.5 Simulations
@@
We utilized open-source MATLAB functions provided by Yeung et al.[30, 31] to generate the simulated EEG data. The simulated data consists of two key components: The signal component is generated to mimic the power spectrum of a typical human EEG recording. The peak component is parameterized to describe the position of the centre of the peak or ERP, its frequency, and its amplitude. These parameters enable us to create sample ERPs, which serve as the basis for testing the effectiveness of our method. The EEG simulation functions provide a setup with 31 electrodes, each sampled at a frequency of 200Hz, with an epoch duration of 0.8 seconds. To generate independent samples, we averaged the random signals over varying numbers of trials, resulting in single 31x200-dimensional samples that closely resemble real EEG data.
For our experimental setup, we aimed to replicate conditions akin to typical comparisons between participants in time-locked Visual Short-Term Memory (VSTM) tasks. We created 20 independent samples consisting solely of EEG time-series, aligning with the EEG power spectrum of a typical human. In parallel, we generated 20 independent samples with specific ERPs, including the N100 and P300 components. The Amplitude of the General EEG signal was set at 10.
@@
The N100 component was parameterized with an amplitude of -5 (typically negative) and a frequency of 15 Hz, with a centre position at 25 frames (around 100ms) considering the 0.8-second epoch. The P300 component was parameterized with an amplitude of 5 (positive and larger than N100) and a frequency of 5 Hz, with a centre position at 75 frames (around 300ms) within the 0.8-second epoch. To introduce realistic variability, the functions incorporate temporal jitter at the onset of the ERPs, mirroring the kind of activity observed in actual EEG ERP data. We then added random Gaussian white noise to the samples containing the simulated ERPs to test the robustness of our FAST Filtering method in the presence of variable levels of external noise. This approach allows us to rigorously test the performance of our method in distinguishing between the presence and absence of these specific ERP components in simulated EEG signals.
The N100 has been previously implicated in various neurological disorders such as Schizophrenia and Attention Deficit Hyperactivity Disorder (ADHD)[32]. The P300 is characterized by a positive deflection in the EEG signal and usually occurs around 300 milliseconds post the presentation of stimuli. The P300 can be influenced by the given task the participant is involved in and is associated with the evaluation of the relevance of stimuli. The P300 has been heavily researched in AD [33, 35, 34] and is associated with decision-making and working memory.
2.6 Visual short-term memory data
This study examines patients with mild cognitive impairment (MCI) due to Alzheimer’s disease (AD), focusing on the pre-dementia stage. The MCI participants are categorized into familial (MCI-FAM) and sporadic (MCI-SPO) groups. Familial AD participants exhibit AD symptoms but do not yet meet the criteria for dementia, although they will inevitably progress to it. Sporadic AD participants, representing the most common type, have an undetermined risk of developing dementia. Both groups are compared to control participants without genetic mutations and free of psychiatric or neurological disorders.
All participants gave written informed consent following the Helsinki Declaration. The Ethics Committees of the Institute of Cognitive Neurology (INECO) and the University of Antioquia approved the study.
2.6.1 Sporadic Mild Cognitive Impairment
Table 1 examines sporadic mild cognitive impairment (MCI-SPO) focusing on demographic and clinical characteristics of the subjects.
Table 1: Demographic and Clinical Characteristics of Subjects MCI-SPO
MCI Patients (n=13) Healthy Controls (n=19)
Age (years) $73.08\pm 9.01$ $67.21\pm 10.14$
Education (years) $14.08\pm 4.44$ $16.50\pm 1.99$
MMSE Scores $26.46\pm 2.47$ $29.50\pm 0.52$
We report on the MMSE, with a detailed clinical and neuro-psychological profile available in Pietto et al. (2016). Patients exhibited multiple-domain amnestic MCI based on various tests. Nine patients were at high risk for AD conversion, while three had non-amnestic MCI multi-domain. The data include 128-channel EEG activity recorded at 512 Hz using a Biosemi Active Two System, filtered from 1 to 100 Hz, and down-sampled to 256 Hz.
2.6.2 Familial Alzheimer’s Disease Dataset
The MCI-FAM group carries the E280A mutation of the presenilin-1 gene, leading to guaranteed early-onset familial AD. Table 2 details the basic Demographic and Clinical Charecterstics as with MCI-SPO.
Table 2: Demographic and Clinical Characteristics of Subjects MCI-FAM
Patients (n=10) Healthy Controls (n=10)
Age (years) $44.4\pm 3.2$ $44.3\pm 5.6$
Education (years) $7.3\pm 4.1$ $6.8\pm 2.9$
MMSE Scores $25.20\pm 4.50$ $29.10\pm 1.10$
The data consist of 60-channel EEG activity recorded with a 64-channel EEG cap using SynAmps 2.5 in Neuroscan at 500 Hz, band-pass filtered from 1 to 100 Hz with impedances below 10 k. Four ocular channels were discarded after being used to factor out oculomotor artifacts.
@@ 2.6.2 Familial Alzheimer’s Disease Dataset
2.6.3 Visual Short-Term Memory Binding and Shape Task Description and Performance
@@
Response accuracy to the two VSTM task conditions was similar for both controls (Mann-Whitney U: 34, Z = 1.17, p = .24, d = .64) and MCI-FAD patients (Mann-Whitney U: 28, Z = 1.63, p = .10, d = .77). However, between-group comparisons showed that controls had higher accuracy in the shape-color binding condition (Mann-Whitney U: 22.5, Z = -2.08, p $<$ .05, d = .93), with no significant differences observed in the shape-only condition (Mann-Whitney U: 25.0, Z = -1.89, p = .063, d = 1.02).
Within-group comparisons showed no significant differences in response accuracy between task conditions for controls (Mann-Whitney U: 66.5, Z = 1.42, p = .16, d = .60) or MCI-SPO patients (Mann-Whitney U: 54, Z = 1.54, p = .12, d = .64). MCI-SPO patients performed significantly worse than controls in both the shape-only (Mann-Whitney U: 42.5, Z = 2.33, p $<$ .05, d = .91) and shape-color binding (Mann-Whitney U: 42.0, Z = 2.35, p $<$ .05, d = .92) conditions.
@@
Signal pre-processing was performed to get signals band-passed into Delta (0.01-4Hz), Theta (4-8Hz), Alpha (8-12Hz), Beta (12-30Hz) and Gamma ($>$30Hz) frequencies with each epoch lasting 1 second post-stimuli exposure.
2.7 Statistical Methods
We designate as a result of clinical interest as a given temporal window where there is a significant difference between patients and controls in the binding task and no significant difference in the shape task as this points to a specific binding deficit in AD.
@@ 2.7 Statistical Methods
The network analysis was done using FAST Connectivity. We first applied Definition 1 to create our FAST filter of overall general VSTM task activity. A separate filter was created for the MCI-FAM dataset and for the MCI-SPO dataset due to them having different experimental parameters (i.e number of electrodes). We then split the FAST Connectivity tensors into 10 0.1 second, non-overlapping temporal windows by averaging over smaller time windows to give us 10 matrices of FAST connectivity for each participant with a high temporal resolution.
@@
Non-parametric Wilcox-on rank-sum test are performed to assess for statistical significance between patients and controls. These are computed at each 0.1s temporal window between the vectors of mean network metrics for patients and controls at each temporal window. This is repeated for the mean edge weights and the mean weighted clustering coefficient values.The direction and size of the differences are calculated using Cohen’s d effect size. In our experiments, a negative value would indicate a greater Local Dirichlet energy in the AD patients for the given VSTM task.
@@
In order to account for the multiple comparisons we applied the Benjamini-Hochberg False Discovery Rate (FDR) correction to account for multiple temporal significance testing. This was done at the 10 and 5 percent level. While 5 percent is often held as the strict standard, the 10 percent level allows us to look for sensitivity pointing towards repeatability across datasets– i.e. where one dataset passed FDR at a given time point at 5 percent and the other at 10 percent.
3 Results
3.1 FAST Connectivity Outperforms Wavelet Transform and Unfiltered Connectivity at Robust,Temporally precise ERP detection
@@ 3 Results
A simulation in a general case scenario is implemented to show the effectiveness of our method at picking up relevant Event-Related Potentials. This is tested across varying noise levels and the number of task trials. First, we performed a three-way comparison between the unfiltered node-function, individual GVD filters and our FAST filter. Our aim is to evaluate the effectiveness of our methodology in detecting these simulated ERPs compared to a ‘control’ group where the ERPs are not present.
Gaussian white noise is added randomly to each of the electrodes equally in the simulated setup (Note, we are adding random white noise which is distinct from the signal generated by the MATLAB functions that resemble the power spectrum of a typical human EEG recording).
Figure 1: A visual demonstration of the effect of the FAST filter on the EEG instantaneous connectivity profile. The Left panel shows the raw EEG connectivity profile, while the Right panel displays the connectivity profile after applying the FAST filter, illustrating the filter’s ability to enhance relevant connectivity patterns and reduce noise.
Figure 1 shows visually the effect of the FAST filter on instantaneous EEG connectivity profiles. We used a time-step consisting of a simulated ERP in the presence of a high level of Gaussian distributed noise. We can see the FAST filter has first established the highly important connections in terms of long-term stable connectivity and identified regions of lower consistency where spurious connections are likely to be present. After the application of the FAST filter we can see that areas of importance have been emphasized (Red patches) while spurious connections and noise has been weighted down. Note that the FAST filter is task-selective and that only areas that are both instantaneously and globally consistent are emphasized. I.e if we had a very weak connection instantaneously but it was strong globally it would not be emphasized. This is a justification for us using both the Shape and Binding tasks to compute our FAST filter as the FAST filter would activate binding specific areas and shape specific areas dependant on the task being analyzed while common strongly connected areas in both tasks would be automatically emphasized. This allows us to utilize the stable connectivity information of both tasks while making sure we do not obtain spurious results due to differences in the tasks. While we can do this due to the similarity of the VSTM binding and shape task and given that they are on the same time-scale, highly differing tasks on different time-scales would likely not be suitable to construct a FAST filter on as the likelihood of spurious results would increase.
Although GVSA is usually performed with distinct individual filters of long-term stable connectivity distinct for each participant, this would result in individual variation and noise in the EEG signals to bias the instantaneous connectivity profiles resulting in False Positives or Type 1 errors.
(a) Individual Filters
(b) FAST Filtered
Figure 2: P-value plots over time for the detection of the N100 and P300 ERP’s (Individual Filters vs FAST filter)
Figure 2 illustrates this, when we use individual filters the noise in the medium is resulting in a very high number of Type 1 errors, with spurious significant difference being picked out (All time-steps are significant with the mean clustering coefficient and edge weights overlapping). FAST connectivity , as shown above, results in no False Positives.
Common baseline EEG analyses revolves around analysis of the power spectrum of the frequency domain of EEG signals. Therefore, it is important to study the added benefits of our approach for classification of EEG signals beyond power spectral analysis. In this vein, we compare FAST functional connectivity with the Wavelet Transform, a common approach to analyzing temporally varying signal power spectra. The CWT produces wavelet coefficients, which capture how the signal correlates with the wavelet at different scales (frequencies) and translations (times). The wavelet coefficients obtained from the CWT are complex numbers. The magnitude squared of these coefficients represents the power of the signal at different frequencies and times.
We compute the overall power of the signal within each temporal window across all channels. We then computed this for all simulated participants or samples with and without a simulated ERP and performed the rank-sum test for significance with FDR correction at varying levels of trials and external noise levels.
@@
We decided to test FAST connectivity’s robustness to noise more rigorously by varying levels of added Gaussian noise at different numbers of trials for each simulated participant EEG. We extracted the FDR corrected $p$-values at the P300 ERP time steps of interest (pre-determined to exist at these times-steps) enabling us to compare the ability of unfiltered dynamic connectivity (the squared difference node function with a support of 1’s with 0’s on the main diagonal as the ’filter’) against the FAST filtered approach to pick out significant differences at these points while also making a comparison with the more traditional wavelet power spectra approach. We tried trials ranging from 50-300 corresponding to typical real-life experiments where EEG signals are recorded.
Figures 3(a), 3(b) and 3(d) shows the $p$-values obtained using unfiltered functional connectivity, FAST functional connectivity, and Wavelet analysis, respectively. Our FAST functional connectivity approach consistently and strongly outperforms the baseline methods in almost all cases. The Wavelet Power spectra analysis, at first glance, performs adequately being able to detect the ERP at all trial sizes when no external noise is added, on further inspection looking at the trend of p-values, the lowest p-value comes at 50 trials, this is unexpected behaviour and the expected pattern emerges at 100 trials and above. This is a strong indication of a Type 1 error taking place due to the inherent noise in the EEG signal causing spurious ’significant differences’. Upon the further observation of the p-values of the Wavelet Transform at all time-steps we noticed that while the ERP could be detected at an external noise level of 0 there was a large number of False Positives at other non-ERP time-steps reducing the power of our statistical analysis greatly. Referring back to Figure 2, this is similar behaviour when the Individual Filters of GVSA is used. Importantly, we noticed that FAST Connectivity had a very low (almost 0) false positive rate, with significant differences only at the time-steps of interest.
(a) Unfiltered
(b) FAST Filtered
(c) Temporal Precision
(d) Wavelet Transform
@@
Figure 3: The $p$-values at a pre-determined time-step of the simulated P300 at increasing levels of trial size and external Gaussian noise. (a) Unfiltered, (b) FAST Filtered, (c) Temporal precision after applying the FAST filter, and (d) Wavelet Spectral analysis.
@@
Another important consideration in analyzing Dynamic functional connectivity in small temporal windows is the temporal resolution we can achieve while still maintaining robustness to noise. Previous methods depended heavily on the length of the sliding window in finding a trade-off between temporal resolution and robustness to noise [25]. We decided to test the window length dependency of FAST connectivity by setting the number of windows equal to the sampling rate, i.e. maximum temporal resolution and repeating our varying trials and external noise level analysis.
In Figure 3(c), we can see the FAST filter’s robustness to increasing levels of noise with the edge weights picking up significant results at the correct time steps in almost all cases in the 100-250 trial range. At 50 trials, it still performs relatively well but there is a decreasing performance at very high levels of external noise is added. The trend shows that as the number of trials increases the detection ability improves while increasing noise decreases this detection ability. The mean edge weights in the unfiltered case fail to detect these simulated ERPs in all levels of noise and trial sizes.
@@
The findings when we set the number of windows to the sampling rate exhibit promise, revealing that despite a reduction in performance when utilizing maximum temporal resolution with 200 windows in contrast to 10 temporal windows as there is clear failure at higher levels of noise at 50 trials (although the ERP is detected at 50 trials at no external noise (p-value 0.04941), the decline is minimal particularly with an optimal number of trials. This observation underscores the robustness of FAST connectivity in variance to window length variations, highlighting its comparative advantage over existing methodologies in capturing dynamic functional connectivity changes.
@@
3.2 Application to visual short-term memory binding in Alzheimer’s disease